OpenStudy (anonymous):

What is the vertex for the function? y = x2 - 4x - 3 (-1, 2) (0, -3) (1, -6) (2, -7 @jim_thompson

jimthompson5910 (jim_thompson5910):

you did this one last night?

OpenStudy (anonymous):

I know how to get one of the vertex which is x=4 well how do I find y

jimthompson5910 (jim_thompson5910):

it's not x = 4

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -(-4)/(2*1) x = 4/2 x = 2

jimthompson5910 (jim_thompson5910):

the x coordinate of the vertex is x = 2 plug it into the function to find y

OpenStudy (zale101):

Follow the steps Jim did and find the function of y.

OpenStudy (anonymous):

how did you get 2?

jimthompson5910 (jim_thompson5910):

in this case, b = -4 and a = 1 you then plug those into x =-b/(2a) to get x = 2

jimthompson5910 (jim_thompson5910):

remember that y = x^2 - 4x - 3 is the same as y = 1x^2 - 4x - 3

OpenStudy (anonymous):

o okay thank you!

jimthompson5910 (jim_thompson5910):

sure thing